Graph the inequality 3x +2y 530
Answer #1

3x +2y <30 3x +2y = 30 If x=0 360) +2y = 30 Y = 30 y = is units. If Y=0 3x +260) = 30 3 X to = 30 = 30 3 x = 10 IIS This ineq

3x+2y < 30 0 2 4 6 F 10 12 14 16